# onlineReferences

### v5

@@ A.1   Non-binary, ordered treatment
=$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
@@
This follows from Assumption 4 and 5.
@@ A.2.1 Two time periods
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1.$
@@
By similar arguments in the proof of Theorem 1 (Theorem 4), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period $1$ under Assumptions 14-19 (thus, we omit the proof for brevity).
@@ A.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 8 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In multiple periods and cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
@@
When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 14, and the following parallel trends assumptions and relevance condition.
Assumption 20 (Parallel trends assumption in the treatment in multiple time periods).
$\text{For all}\hskip 2.84526pts<t,\hskip 2.84526ptE[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=s]\hskip 2.84526pt\text{is identical for all}\hskip 2.84526pte\in\mathcal{S}(E_{i}).$
Assumption 21 (Parallel trends assumption in the outcome in multiple time periods).
For $\text{For all}\hskip 5.69054pts<t,\hskip 5.69054ptE[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=s]$
is identical for $\text{is identical for all}\hskip 5.69054pte\in\mathcal{S}(E_{i}).$
Assumption 22 (Relevance condition in multiple time periods).
$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i}-D^{\infty}_{i}|E_{i}=e,T_{i}=e+l]>0.$
A.3   Introducing the treatment path in $2\times 2$ DID-IV designs
We discuss the DID-IV identification strategy in two periods and two groups with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. In this setting, our target parameter is the LATET’ defined below.
Definition.
The local average treatment effect on the treated with the treatment path (LATET’) is
$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
where $\Delta_{0}$ and $\Delta_{1}$ are:
$\Delta_{0}\equiv E[Y_{1}((1,1))-Y_{1}((1,0))|E_{i}=1,(CM^{Z}\land DF^{T})],$
$\Delta_{1}\equiv E[Y_{1}((0,1))-Y_{1}((0,0))|E_{i}=1,(CM^{Z}\land NT^{T})].$
The weights $w_{0}$ and $w_{1}$ are:
$w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
We consider the identifying assumptions for the Wald-DID estimand to identify the LATET’. First, we make the following no anticipation assumptions on the outcome instead of imposing Assumption 1 (No carryover assumption).
Assumption 23 (No anticipation of the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
$\forall z\in\mathcal{S}(Z),Y_{i,0}((0,1),z)=Y_{i,0}((0,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land NT^{T}).$
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$.
Assumption 24 (No anticipation of the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
$\forall z\in\mathcal{S}(Z),Y_{i,0}((1,1),z)=Y_{i,0}((1,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land DF^{T}).$
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$.
We can rationalize Assumptions 23-24 by considering the similarity between these assumptions and the no anticipation assumption made in the recent DID literature; all of them restrict the anticipatory behavior for those who belong to an exposed group (treatment group) and are affected by a policy shock (the treatment in DID designs).
Next, we make the parallel trends assumption in the outcome suitable for the set up considered here. We first define the similar notation as in Subsection 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
$\bar{\Delta}_{AT^{T},e}\equiv E[Y_{i,1}((1,1))-Y_{i,0}((1,1))|AT^{T},E_{i}=e],$
$\bar{\Delta}_{NT^{T},e}\equiv E[Y_{i,1}((0,0))-Y_{i,0}((0,0))|NT^{T},E_{i}=e],$
$\bar{\Delta}_{CM^{T},e}\equiv E[Y_{i,1}((0,1))-Y_{i,0}((0,1))|CM^{T},E_{i}=e],$
$\bar{\Delta}_{DF^{T},e}\equiv E[Y_{i,1}((1,0))-Y_{i,0}((1,0))|DF^{T},E_{i}=e].$
Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups, respectively:
$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},\hskip 5.69054pt\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
where the definition of the weight $w_{g^{t},e}$ is same as in Subsection 2.4.
As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the same expected time gain between exposed and unexposed groups. The following assumption makes such a restriction in the current set up.
Assumption 25 (Parallel trend assumption in the outcome).
The expected time gain between exposed and unexposed groups is the same: $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
The theorem below shows that the Wald-DID estimand is equal to the LATET’ under Assumptions 2-5, 7 and Assumptions 23-25.
Theorem 5.
Suppose Assumptions 2-5, 7 and Assumptions 23-25 hold. Then, the Wald-DID estimand corresponds to the LATET’.
Proof.
First, we consider the denominator of the Wald-DID estimand. From the argument in the proof of Theorem 1, one can show the following result:
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]=Pr(CM^{Z}|E=1)$
$=Pr(CM^{Z}\land DF^{T}|E=1)+Pr(CM^{Z}\land NT^{T}|E=1).$ (15)
Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
$E[Y_{1}-Y_{0}|E=0]=\bar{\Delta}_{0}.$ (16)
This follows from Table 2 and the definition of the expected time gain $\bar{\Delta}_{0}$.
Finally, we decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
$E[Y_{1}-Y_{0}|E=1]=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{0}((1,1))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+\bar{\Delta}_{NT^{T},1}w_{NT^{T},1}+\bar{\Delta}_{DF^{T},1}w_{DF^{T},1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$=\bar{\Delta}_{1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$ (17)
The first equality follows from Table 1 and Assumption 2. The second equality follows from Assumption 23, Assumption 24 and simple manipulation. The third equality follows from Table 1. The final equality follows from the definition of the expected time gain $\bar{\Delta}_{1}$.
From the result (16) with (17) and Assumption 25, we obtain
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1).$ (18)
Combining the result (18) with (15), we obtain the desired result. ∎
@@ Appendix B Deriving the influence function in repeated cross section settings
[…]
@@
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Using similar arguments in the proof of Theorem 3, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 14, 20-22, and have the following influence function:
@@
×$\times\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg{)},$ (19)

### v6

@@ A.1  Non-binary, ordered treatment
$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
@@
This follows from Assumptions 45.
@@ A.2.1 Two time periods
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
Here, we consider the monotonicity assumption that rules out the defiers, as our target parameter measures the treatment effects among the compliers.
@@
By similar arguments in the proof of Theorem 1 (Theorem 5), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period 11 under Assumptions 1924 (thus, we omit the proof for brevity).
@@ A.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 8 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 19 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}=\min\{t:Z_{i,t}=1\}\in\{2,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $E_{i,e}=\mathbf{1}\{E_{i}=e\}$ denote the binary indicator that takes one if unit $i$ belongs to cohort $e$. Let $\bar{e}=\max_{i=1,\dots,n}E_{i}$ denote the largest cohort value in the dataset. Let $\mathcal{E}=\mathcal{S}(E_{i})\setminus\{\bar{e}\}\subseteq\{2,3,\dots,T\}$ denote the support of $E_{i}$ excluding $\bar{e}$.
Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In multiple periods and cohorts settings with a binary treatment, our target parameter is the $CLATT_{e,e+l}$ in period $T=e+l$:
@@
If we have an ordered, non-binary treatment, our target parameter is the $CACRT_{e,e+l}$ in period $T=e+l$:
$CACRT_{e,e+l}\equiv\sum_{j=1}^{J}w^{e}_{e+l,j}\cdot E[Y_{i,e+l}(j)-Y_{i,e+l}(j-1)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}]$
where the weight $w^{e}_{e+l,j}$ is:
$w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|T_{i}=e+l,E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|T_{i}=e+l,E_{i}=e)}.$
Similar to Section 3.3, we first consider the following estimand to identify each $CLATT_{e,e+l}$:
$w^{DID,r}_{e,l}=\frac{\Delta^{l}_{Y,E=e}-\Delta^{l}_{Y,E=\infty}}{\Delta^{l}_{D,E=e}-\Delta^{l}_{D,E=\infty}},\qquad\text{for }e\in\mathcal{E}\text{ and }l\in\{0,\dots,T-e\}.$
where, for any $e\in\mathcal{S}(E_{i})$,
$\Delta^{l}_{Y,E=e}$ $=E[Y_{i}\mid E_{i}=e,T_{i}=e+l]-E[Y_{i}\mid E_{i}=e,T_{i}=e-1],$
$\Delta^{l}_{D,E=e}$ $=E[D_{i}\mid E_{i}=e,T_{i}=e+l]-E[D_{i}\mid E_{i}=e,T_{i}=e-1].$
Note that in this Wald-DID estimand, the pre-exposed period is e1e-1 and the control group is the never exposed cohort. Here, we assume that the largest cohort value in the dataset is $\bar{e}=\infty$.
@@
When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 19, and the following parallel trends assumptions and the relevance condition.
Assumption 25 (Parallel trends assumption in the treatment based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=t-1]$
$E[D_{i}^{\infty}|E_{i}=\infty,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=\infty,T_{i}=t-1].$
Assumption 26 (Parallel trends assumption in the outcome based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t-1]$
$E[Y_{i}(D_{i}^{\infty})|E_{i}=\infty,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=\infty,T_{i}=t-1].$
Assumption 27 (Relevance condition based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\},$
$\Delta^{l}_{D,E=e}-\Delta^{l}_{D,E=\infty}>0.$
By similar arguments in the proof of Theorem 2, one can show that each Wald-DID estimand $w^{DID,r}_{e,l}$ is equal to the $CLATT_{e,e+l}$ ($CACRT_{e,e+l}$ under an ordered, non-binary treatment) for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ under Assumptions 8, 11, 12 in Section 3, Assumption 19, and Assumptions 2527 (thus, we omit the proof for brevity).
When there is no never-exposed cohort, or its sample size is too small, one can instead consider the following estimand:
$w^{DID,ny,r}_{e,l}=\frac{\Delta^{l}_{Y,E=e}-\Delta^{l}_{Y,ny}}{\Delta^{l}_{D,E=e}-\Delta^{l}_{D,ny}},$
where
$\Delta^{l}_{Y,ny}$ $=E[Y_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e+l]-E[Y_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e-1],$
$\Delta^{l}_{D,ny}$ $=E[D_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e+l]-E[D_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e-1].$
for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. In this Wald-DID estimand, the control group is the not-yet-exposed cohorts.
If we adopt the above estimand $w^{DID,ny,r}_{e,l}$, we can replace Assumptions 2527 with Assumptions 2830 below.
Assumption 28 (Parallel trends assumption in the treatment based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=t-1]$
$E[D_{i}^{\infty}|Z_{i}=0,E_{i,e}=0,T_{i}=t]-E[D_{i}^{\infty}|Z_{i}=0,E_{i,e}=0,T_{i}=t-1].$
Assumption 29 (Parallel trends assumption in the outcome based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t-1]$
$E[Y_{i}(D_{i}^{\infty})|Z_{i}=0,E_{i,e}=0,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|Z_{i}=0,E_{i,e}=0,T_{i}=t-1].$
Assumption 30 (Relevance condition based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptl<\bar{e}-e,$
$\Delta^{l}_{D,E=e}-\Delta^{l}_{D,ny}>0.$
Again, by similar arguments in the proof of Theorem 3, one can show that each Wald-DID estimand $w^{DID,ny,r}_{e,l}$ is equal to the $CLATT_{e,e+l}$ ($CACRT_{e,e+l}$ under an ordered, non-binary treatment) for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$ under Assumptions 8, 11, 12 in Section 3), 19, and Assumptions 2830 (thus, we omit the proof for brevity).
@@ Appendix B Deriving the influence function in repeated cross section settings
In this appendix, we present the influence function of our STS estimator when repeated cross section data are available. [^22] The derivation here is essentially the same as in de Chaisemartin and D’Haultfœuille (2018), who present the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups. Let $\theta_{e,l}^{r}$ define
@@
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Then, if we set $U=\{\infty\}$ ($U=\{e^{\prime}\in\mathcal{S}(E_{i}):e^{\prime}>e+l\}$), one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 19 and Assumptions 2527 (Assumptions 2830), and have the following influence function:
@@
$\times\Bigg(\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg),$ (14)